# Sprecher Networks: A Parameter-Efficient Kolmogorov-Arnold Architecture

### v1

@@ Abstract
We present Sprecher Networks (SNs), a family of trainable neural architectures inspired by the classical Kolmogorov–Arnold–Sprecher (KAS) construction for approximating multivariate continuous functions. Distinct from Multi-Layer Perceptrons (MLPs) with fixed node activations and Kolmogorov-Arnold Networks (KANs) featuring learnable edge activations, SNs utilize shared, learnable splines (monotonic and general) within structured blocks incorporating explicit shift parameters and mixing weights. Our approach directly realizes Sprecher’s specific 1965 “sum of shifted splines” formula in its single-layer variant and extends it to deeper, multi-layer compositions. We further enhance the architecture with optional lateral mixing connections that enable intra-block communication between output dimensions, providing a parameter-efficient alternative to full attention mechanisms. Beyond parameter efficiency with $O(LN+LG)$ scaling (where $G$ is the knot count of the shared splines) versus MLPs’ $O(LN^{2})$, SNs admit a sequential evaluation strategy that reduces peak forward-intermediate memory from $O(N^{2})$ to $O(N)$ (treating batch size as constant), making much wider architectures feasible under memory constraints. We demonstrate empirically that composing these blocks into deep networks leads to highly parameter and memory-efficient models, discuss theoretical motivations, and compare SNs with related architectures (MLPs, KANs, and networks with learnable node activations).
@@ 1 Introduction and historical background
Approximation of continuous functions by sums of univariate functions has been a recurring theme in mathematical analysis and neural networks. The Kolmogorov–Arnold Representation Theorem [10, 1] established that any multivariate continuous function $f:[0,1]^{d}\to\mathbb{R}$ can be represented as a finite composition of continuous functions of a single variable and the addition operation. Specifically, Kolmogorov (1957) showed that such functions can be represented as a finite sum involving univariate functions applied to sums of other univariate functions of the inputs.
@@
for a single monotonic inner function $\phi$, a continuous outer function $\Phi$, a constant shift parameter $\eta>0$, and constants $\lambda_{p}$. This construction simplified the representation by using only one inner function $\phi$, relying on shifts of the input coordinates ($x_{p}+\eta q$) and an outer summation index shift (+q+q) to achieve universality. The key insight of shifting input coordinates and summing evaluations under inner and outer univariate maps is central to Sprecher’s specific result.
@@
Modern context. Understanding how novel architectures relate to established ones is crucial. Standard Multi-Layer Perceptrons (MLPs) [7] employ fixed nonlinear activation functions at nodes and learnable linear weights on edges, justified by the Universal Approximation Theorem [4, 8]. Extensions include networks with learnable activations on nodes, sometimes called Adaptive-MLPs or Learnable Activation Networks (LANs) [6, 17, 13], which retain linear edge weights but make the node non-linearity trainable. Recent work has revitalized interest in leveraging Kolmogorov-Arnold representations for modern deep learning. Notably, Kolmogorov-Arnold Networks (KANs) [13] were introduced, proposing an architecture with learnable activation functions (splines) placed on the edges of the network graph, replacing traditional linear weights and fixed node activations. This enables KANs to encode structure that MLPs must discover implicitly, leading to benefits such as better scalability, and improved performance on physics-based tasks. Due to their success, KANs have been extended in various forms, and adapted to a variety of tasks – see [14] for a survey. To name a few, alternative learned activations have been explored [12, 3], and several architectural extensions have been developed, e.g. convolutional [2], recurrent [5], and attention-based [16]. Sprecher Networks (SNs), as we detail below, propose a distinct approach derived directly from Sprecher’s 1965 formula. SNs employ function blocks containing shared learnable splines ($\phi,\Phi$), learnable mixing weights ($\lambda$), explicit structural shifts ($\eta,q$), and optionally, lateral mixing connections for intra-block communication. This structure offers a different alternative within the landscape of function approximation networks. Code is available at https://github.com/Zelaron/Sprecher-Network.
@@ 2 Motivation and overview of Sprecher Networks
While MLPs are the workhorse of deep learning, architectures inspired by KAS representations offer potential benefits, particularly in interpretability and potentially parameter efficiency for certain function classes. KANs explore one direction by placing learnable functions on edges. Our Sprecher Networks (SNs) explore a different direction, aiming to directly implement Sprecher’s constructive formula within a trainable framework and extend it to deeper architectures.
@@
  • Each functional block (mapping between layers) is organized around a shared monotonic spline $\phi(\cdot)$ and a shared general spline $\Phi(\cdot)$, both learnable.
@@
  • Each block includes learnable mixing weights $\lambda_{i}$ (a vector, not a matrix) that combine contributions from different input dimensions, with the same weights shared across all output dimensions.
@@
Concretely, each Sprecher block applies the transformation:
$(x_{i})_{i=1}^{d_{\mathrm{in}}}\;\mapsto\;\Bigl[\Phi^{(\ell)}\Bigl(s_{q}^{(\ell)}\;+\;\tau^{(\ell)}\!\!\sum_{j\in\mathcal{N}(q)}\omega_{q,j}^{(\ell)}\,s_{j}^{(\ell)}\Bigr)\Bigr]_{q=0}^{d_{\mathrm{out}}-1},$
where $s_{q}^{(\ell)}=\sum_{i=1}^{d_{\mathrm{in}}}\lambda^{(\ell)}_{i}\,\phi^{(\ell)}(x_{i}+\eta^{(\ell)}q)+\alpha q$. Lateral mixing (if enabled) modifies the pre-activation $s_{q}^{(\ell)}$ before $\Phi^{(\ell)}$, whereas residual connections (if enabled) are applied after $\Phi^{(\ell)}$ and do not introduce lateral coupling across $q$. For scalar outputs, the outputs of the final Sprecher block are aggregated (via summation); for vector outputs, no final summation is applied.
@@
Throughout this paper, we denote Sprecher Network architectures using arrow notation of the form $d_{\mathrm{in}}\to[d_{1},d_{2},\ldots,d_{L}]\to d_{\mathrm{out}}$, where $d_{\mathrm{in}}$ is the input dimension, $d_{1},d_{2},\ldots,d_{L}]$ represents the hidden layer dimensions (widths), and $d_{\mathrm{out}}$ is the final output dimension of the network. For scalar output ($d_{\mathrm{out}}=1$), the final block’s outputs are summed. For vector output ($d_{\mathrm{out}}>1$), an additional non-summed block maps from $d_{L}$ to $d_{\mathrm{out}}$. For example, $2\to[5,3,8]\to 1$ describes a network with 2-dimensional input, three hidden layers of widths 5, 3, and 8 respectively, and a scalar output (implying the final block’s outputs of dimension 8 are summed). $2\to[5,3]\to 4$ describes a network with 2-dimensional input, two hidden layers of widths 5 and 3, and a 4-dimensional vector output (implying an additional output block maps from dimension 3 to 4 without summation). When input or output dimensions are clear from context, we may use the abbreviated notation $d_{1},d_{2},\ldots,d_{L}]$ to focus on the hidden layer structure.
@@ 3.1 Sprecher block structure
  • Monotonic spline $\phi(\cdot)$: A non-decreasing spline mapping $\mathbb{R}\to[0,1]$ (piecewise-linear in our implementation; see Section 7) that is applied coordinate-wise. While we denote the domain as $\mathbb{R}$ for generality, in practice each $\phi^{(\ell)}$ operates on a bounded domain determined dynamically by the network’s structure and current parameters, as detailed in Section 7.4. This function is shared across all input-output connections within the block and its coefficients are learnable. Monotonicity is enforced during training through appropriate parameterization (see Section 7).
  • General spline $\Phi(\cdot)$: A spline mapping $\mathbb{R}\to\mathbb{R}$ (piecewise-linear or $C^{1}$ piecewise-cubic Hermite/PCHIP; see Section 7, without monotonicity constraints) whose domain and codomain are determined by the network structure and can be either fixed or made trainable. If trainable codomains are used, they can be parameterized in various ways, such as by center and scale (half-width) parameters, allowing the spline to adapt its output range during training. This function is also shared across the block and its coefficients are learnable.
  • Mixing weights vector $\lambda$: A vector $\{\lambda_{i}\}$ of size $d_{\mathrm{in}}$, whose entries are learnable. These weights linearly combine the contributions from different input dimensions after transformation by $\phi$. Crucially, these weights are shared across all output dimensions within the block, maintaining fidelity to Sprecher’s original formulation.
  • Shift parameter $\eta$: A learnable scalar $\eta$. This parameter controls the magnitude of the input shift $x_{i}+\eta q$, which depends on the output index $q$. While Sprecher’s original construction requires $\eta>0$, practical implementations may allow $\eta$ to take any real value during training.
  • Lateral mixing parameters (optional): When enabled, learnable parameters that allow intra-block communication between output dimensions. This includes a scalar $\tau^{(\ell)}$ (lateral scale) and weight vector(s) $\omega^{(\ell)}$ that enable outputs to exchange information before passing through $\Phi^{(\ell)}$. Two variants are supported: cyclic (each output receives from its cyclic neighbor) and bidirectional (each output mixes with both neighbors).
@@
Concretely, we first define a single Sprecher block as an operator. Let $B^{(\ell)}:\mathbb{R}^{d_{\ell-1}}\to\mathbb{R}^{d_{\ell}}$ denote the $\ell$-th Sprecher block with parameters $\phi^{(\ell)},\Phi^{(\ell)},\eta^{(\ell)},\lambda^{(\ell)}$, a fixed spacing constant $\alpha$ (set to $1$ in all experiments unless noted), and optionally $\tau^{(\ell)},\omega^{(\ell)}$. Given an input vector $\mathbf{x}=(x_{1},\dots,x_{d_{\mathrm{in}}})\in\mathbb{R}^{d_{\mathrm{in}}}$, the block computes its output vector $\mathbf{h}\in\mathbb{R}^{d_{\mathrm{out}}}$ component-wise as:
$B^{(\ell)}(\mathbf{x})]_{q}=\Phi^{(\ell)}\Biggl(\,s_{q}^{(\ell)}+\tau^{(\ell)}\sum_{j\in\mathcal{N}(q)}\omega_{q,j}^{(\ell)}s_{j}^{(\ell)}\Biggr),$
where $s_{q}^{(\ell)}=\sum_{i=1}^{d_{\mathrm{in}}}\lambda_{i}^{(\ell)}\phi^{(\ell)}(x_{i}+\eta^{(\ell)}q)+\alpha q$ represents the pre-mixing activation, $\alpha$ is a fixed channel-spacing constant (we set $\alpha=1$ in all experiments unless explicitly stated otherwise), and $\mathcal{N}(q)$ denotes the neighborhood structure:
@@
  • Cyclic: $\mathcal{N}(q)=\{(q+1)\bmod d_{\mathrm{out}}\}$ with scalar weights $\omega_{q}^{(\ell)}$
  • Bidirectional: $\mathcal{N}(q)=\{(q-1)\bmod d_{\mathrm{out}},(q+1)\bmod d_{\mathrm{out}}\}$ with weights $\omega_{q,\text{bwd}}^{(\ell)}$ for the $(q-1)\bmod d_{\mathrm{out}}$ neighbor and $\omega_{q,\text{fwd}}^{(\ell)}$ for the $(q+1)\bmod d_{\mathrm{out}}$ neighbor
While $\alpha=1$ maintains theoretical fidelity, alternative values may be explored to improve optimization dynamics in deeper networks. Note that $q$ serves dual roles here: as an output index ($q=0,\ldots,d_{\mathrm{out}}-1$) and as an additive shift parameter within the formula.
@@
Figure 2: Internal structure of a Sprecher block showing spline sharing. Unlike KANs where each edge has a unique learnable spline, a Sprecher block uses only two shared splines: one monotonic $\phi$ applied to all shifted inputs, and one general $\Phi$ applied to all weighted sums. The mixing weights $\lambda_{i}$ are shared across all output dimensions. Diversity across outputs arises from the index-dependent shifts $\eta q$ (applied before $\phi$) and $\alpha q$ (added to each sum). This extreme parameter sharing yields $O(d_{\mathrm{in}}+G)$ parameters per block versus $O(d_{\mathrm{in}}\cdot d_{\mathrm{out}}\cdot G)$ for KANs.

### v2

@@ Abstract
We introduce Sprecher Networks (SNs), a family of trainable architectures derived from David Sprecher’s 1965 constructive form of the Kolmogorov–Arnold representation. Each SN block implements a “sum of shifted univariate functions” using only two shared learnable splines per block, a monotone inner spline $\phi$ and a general outer spline $\Phi$, together with a learnable shift parameter $\eta$ and a mixing vector $\lambda$ shared across all output dimensions. Stacking these blocks yields deep, compositional models; for vector-valued outputs we append an additional non-summed output block.
We also propose an optional lateral mixing operator enabling intra-block communication between output channels with only $O(d_{\mathrm{out}})$ additional parameters. Owing to the vector (not matrix) mixing weights and spline sharing, SNs scale linearly in width, approximately $O\!\left(\sum_{\ell}(d_{\ell-1}+d_{\ell}+G)\right)$ parameters for $G$ spline knots, versus $O\!\left(\sum_{\ell}d_{\ell-1}d_{\ell}\right)$ for dense MLPs and $O\!\left(G\sum_{\ell}d_{\ell-1}d_{\ell}\right)$ for edge-spline KANs. This linear width-scaling is particularly attractive for extremely wide, shallow models, where low depth can translate into low inference latency. Finally, we describe a sequential forward implementation that avoids materializing the $d_{\mathrm{in}}\times d_{\mathrm{out}}$ shifted-input tensor, reducing peak forward-intermediate memory from quadratic to linear in layer width (treating batch size as constant), which is relevant in memory-constrained settings such as on-device/edge inference; we demonstrate deployability via fixed-point real-time digit classification on a resource-constrained embedded device with only 4 MB RAM. We provide empirical demonstrations on supervised regression, Fashion-MNIST classification (including stable training at 25 hidden layers with residual connections and normalization), and a Poisson PINN, together with controlled comparisons to MLP and KAN baselines.
@@ 1 Introduction and historical background
Approximation of continuous functions by sums of univariate functions has been a recurring theme in mathematical analysis and neural networks. The Kolmogorov–Arnold Representation Theorem [10, 1] established that any multivariate continuous function $f:[0,1]^{n}\to\mathbb{R}$ can be represented as a finite composition of continuous functions of a single variable and the addition operation. Specifically, Kolmogorov (1957) showed that such functions can be represented as a finite sum involving univariate functions applied to sums of other univariate functions of the inputs.
@@
for a single monotone inner function $\phi$, a continuous outer function $\Phi$, a constant shift parameter $\eta>0$, and constants $\lambda_{p}$. This construction simplified the representation by using only one inner function $\phi$, relying on shifts of the input coordinates ($x_{p}+\eta q$) and an outer summation index shift (+q+q) to achieve universality. The key insight of shifting input coordinates and summing evaluations under inner and outer univariate maps is central to Sprecher’s specific result.
@@
Modern context. Understanding how novel architectures relate to established ones is crucial. Standard Multi-Layer Perceptrons (MLPs) [7] employ fixed nonlinear activation functions at nodes and learnable linear weights on edges, justified by the Universal Approximation Theorem [4, 8]. Extensions include networks with learnable activations on nodes, sometimes called Adaptive-MLPs or Learnable Activation Networks (LANs) [6, 19, 12], which retain linear edge weights but make the node non-linearity trainable. Recent work has revitalized interest in leveraging Kolmogorov-Arnold representations for modern deep learning. Notably, Kolmogorov-Arnold Networks (KANs) [12] were introduced, proposing an architecture with learnable activation functions (splines) placed on the edges of the network graph, replacing traditional linear weights and fixed node activations. This enables KANs to encode structure that MLPs must discover implicitly, leading to benefits such as better scalability, and improved performance on physics-based tasks. Due to their success, KANs have been extended in various forms, and adapted to a variety of tasks – see [13] for a survey. To name a few, alternative learned activations have been explored [11, 3], and several architectural extensions have been developed, e.g. convolutional [2], recurrent [5], and attention-based [18]. Sprecher Networks (SNs), as we detail below, propose a distinct approach, derived directly from Sprecher’s 1965 formula. SNs employ function blocks containing shared learnable splines ($\phi,\Phi$), learnable mixing weights ($\lambda$), explicit structural shifts ($\eta,q$), and optionally, lateral mixing connections for intra-block communication. This structure offers a different alternative within the landscape of function approximation networks. Code is available at https://github.com/Zelaron/Sprecher-Network.
1.1 Scalability comparison with related architectures
Before detailing the SN architecture, we present a key empirical result that motivates our design choices. Recent works have proposed Kolmogorov-Arnold-inspired architectures claiming “Sprecher-inspired” efficiency, including GS-KAN [15] and SaKAN [16]. While both invoke Sprecher’s 1965 construction as theoretical motivation (and reduce the number of distinct learned univariate functions), they retain dense $O(d_{\mathrm{in}}\times d_{\mathrm{out}})$ weight matrices: GS-KAN uses per-edge learnable weights $\lambda_{p,q}$ connecting all input-output pairs, and SaKAN includes dense residual weight matrices $u_{ij}$ for linear residuals. In contrast, SNs use only vector weights $\lambda_{i}$ (shared across outputs), achieving genuinely $O(d)$ parameter scaling in layer width.
Memory scalability benchmark.
To demonstrate this difference empirically, we conduct a memory scalability stress test (here MLP denotes a stack of dense linear layers with bias): starting at width 512 with depth 3 on 64-dimensional input, we double the width until only one architecture survives without running out of memory (OOM). For each width, we run a single training step (forward, mean-squared error, backward, Adam update) and record the peak additional memory used (relative to a cleaned baseline), taking the maximum of accelerator-allocated memory and process resident set size (RSS). To isolate connectivity scaling rather than spline resolution, we use a small fixed univariate-function parameterization (i.e., we replace the usual spline parameterizations used elsewhere in the paper for this stress test): the KAN baseline uses 5 learnable coefficients per edge, while our GS-KAN and SaKAN baselines (and the SN) use a 1-parameter piecewise-linear nonlinearity (PReLU; $x\mapsto\max(0,x)+a\,\min(0,x)$ with learnable slope $a$) for each required univariate function. For SN, we additionally use the memory-efficient evaluation strategy described in Section 7.3, which iterates over $q$ (optionally in small chunks) to avoid materializing the full $(B\times d_{\mathrm{in}}\times d_{\mathrm{out}})$ shifted-input tensor. Table 1 shows results on a machine with 8GB of system RAM (exact OOM breakpoints are hardware- and implementation-dependent).
Table 1: Scalability benchmark: parameter count and peak additional memory (MB) during a single Adam training step (forward+backward+update) as layer width increases. Configuration: batch_size=32, input_dim=64, depth=3 hidden layers, output_dim=1 (with an explicit final output layer; for SN, this is a final Sprecher block with $d_{\mathrm{out}}=1$); Apple M-series GPU (Metal/MPS), float32. At width 16384, MLP and SaKAN run out of memory; KAN and GS-KAN were skipped after earlier OOMs; SN uses 49,228 parameters and 7.4 MB peak additional memory.
[…]
512 MLP 559,105 8.8 OK
512 KAN 2,787,840 42.7 OK
512 GS-KAN 559,109 138.3 OK
512 SaKAN 559,109 9.0 OK
512 SN 1,612 0.5 OK
1024 MLP 2,166,785 33.6 OK
1024 KAN 10,818,560 184.1 OK
1024 GS-KAN 2,166,789 536.7 OK
1024 SaKAN 2,166,789 34.0 OK
1024 SN 3,148 1.1 OK
2048 MLP 8,527,873 131.1 OK
2048 KAN 42,608,640 650.6 OK
2048 GS-KAN 8,527,877 2,113.3 OK
2048 SaKAN 8,527,877 132.1 OK
2048 SN 6,220 2.0 OK
4096 MLP 33,832,961 518.2 OK
4096 KAN 169,103,360 2,581.6 OK
4096 GS-KAN 33,832,965 OOM
4096 SaKAN 33,832,965 520.1 OK
4096 SN 12,364 4.2 OK
8192 MLP 134,774,785 2,060.4 OK
8192 KAN 673,751,040 OOM
8192 GS-KAN 134,774,789 SKIP(after OOM)
8192 SaKAN 134,774,789 2,064.3 OK
8192 SN 24,652 8.3 OK
16384 MLP 537,985,025 OOM
16384 KAN 2,689,679,360 SKIP(after OOM)
16384 GS-KAN 537,985,029 SKIP(after OOM)
16384 SaKAN 537,985,029 OOM
16384 SN 49,228 7.4 OK
At width 4096, GS-KAN runs out of memory and is therefore not re-run at larger widths (reported as SKIP(after OOM)). In a straightforward fully vectorized evaluation, GS-KAN can OOM early despite MLP-like parameter counts because evaluating the shifted activations $\psi(x_{p}+\epsilon_{q})$ for all $(p,q)$ at once materializes a tensor of shape $B\times N_{\mathrm{in}}\times N_{\mathrm{out}}$ (here $B=32$). At width 8192, standard KAN runs out of memory and is likewise skipped at width 16384. At width 16384, both MLP and SaKAN run out of memory, leaving SN as the only architecture that completes the step (49,228 parameters, 7.4 MB peak additional memory). For reference, the computed parameter counts at this scale are 537,985,025 (MLP), 537,985,029 (SaKAN/GS-KAN), and 2,689,679,360 (KAN).
Capacity utilization.
A natural question is whether the extreme parameter efficiency comes at the cost of learning capacity. To verify that the wide SN actually uses its capacity rather than merely allocating it, we train the width-16384 SN (using the same lightweight configuration as in Table 1, including the memory-efficient sequential evaluation from Section 7.3) on a 64-dimensional regression task adapted from [12]: $f(\mathbf{x})=\exp\bigl(\frac{1}{64}\sum_{i=1}^{64}\sin^{2}(\frac{\pi x_{i}}{2})\bigr)$ for $\mathbf{x}\in[0,1]^{64}$. We optimize mean-squared error on a fixed batch of 32 random samples for 400 epochs with Adam (learning rate $0.001$). Figure 1 shows the best-so-far training loss (the minimum loss observed up to each epoch), which is robust to occasional loss spikes in this high-dimensional setting. The best loss decreases from $3.3867e+00$ to 6.7672e026.7672e-02, indicating that the model remains trainable and can substantially reduce loss even at extreme parameter efficiency.
[…]
Figure 1: Training curve for SN at width 16384 (the only model that completes the memory benchmark at this scale) on a 64-dimensional regression task. The best-so-far training loss decreases from 3.3867e+$0$ to 6.7672e-02 over 400 epochs, indicating that the model remains trainable and can substantially reduce loss despite its extreme parameter efficiency.
Why competitors fail to scale.
The results in Table 1 reflect two distinct scaling bottlenecks: (i) quadratic parameter growth from dense $N_{\mathrm{in}}\times N_{\mathrm{out}}$ matrices (and their optimizer state), and (ii) quadratic edge-activation working memory when edgewise nonlinearities are evaluated for all $(p,q)$ at once. GS-KAN [15] parameterizes edge functions as $\lambda_{p,q}\cdot\psi(x_{p}+\epsilon_{q})$, where $\lambda_{p,q}$ is a learnable weight for each input-output pair (p,q)(p,q)—this retains a dense $d_{\mathrm{in}}\times d_{\mathrm{out}}$ matrix and yields $O(d_{\mathrm{in}}\times d_{\mathrm{out}})$ parameters per layer despite sharing the basis function $\psi$. In a straightforward fully vectorized evaluation, evaluating $\psi(x_{p}+\epsilon_{q})$ for all $(p,q)$ at once materializes a $B\times N_{\mathrm{in}}\times N_{\mathrm{out}}$ activation tensor (here $B=32$), explaining the early OOM of GS-KAN at width 4096 even though its parameter count nearly matches an MLP at the same width. This activation-memory bottleneck is implementation-dependent and can be alleviated with sequential/chunked evaluation, but the dense $\lambda_{p,q}$ matrix still imposes $O(N_{\mathrm{in}}\times N_{\mathrm{out}})$ parameter (and optimizer-state) scaling. SaKAN [16] similarly reduces the number of distinct learned univariate functions but retains a dense linear residual matrix $u_{ij}\in\mathbb{R}^{n_{\mathrm{out}}\times n_{\mathrm{in}}}$ in each layer (see their Eq. 8), so its width-scaling remains dominated by $O(N_{\mathrm{in}}\times N_{\mathrm{out}})$ parameters (even when techniques such as gradient-free spline evaluation reduce spline-gradient memory). Standard KANs [12] place learned functions on every edge, giving $O(G\,N_{\mathrm{in}}\times N_{\mathrm{out}})$ parameters.
In contrast, SNs follow Sprecher’s original construction more faithfully: the mixing weights $\lambda_{i}$ are shared across all output dimensions, collapsing the weight tensor from a full matrix to a single vector and yielding only $O(N_{\mathrm{in}})$ mixing-weight parameters per block (plus the shared spline parameters, independent of $N_{\mathrm{out}}$; optional lateral mixing adds at most linear $O(N_{\mathrm{out}})$ parameters, and cyclic/node residual connections add at most linear $O(\max\{N_{\mathrm{in}},N_{\mathrm{out}}\})$ parameters, whereas a dense linear residual projection would instead contribute $O(N_{\mathrm{in}}N_{\mathrm{out}})$ parameters). Combined with the memory-efficient sequential evaluation described in Section 7.3 (and, when needed, checkpointing when training), this yields linear parameter scaling and small working-memory growth in practice. This linear scaling is particularly advantageous for shallow but very wide networks: while MLPs and KANs quickly exhaust memory as layer width increases (Table 1), SNs can accommodate extremely wide layers with minimal overhead. Such architectures may benefit low-latency settings where fewer sequential blocks reduce depth-induced latency and the per-block computation is highly parallelizable, though the arithmetic cost of a block still scales as $O(d_{\mathrm{in}}d_{\mathrm{out}})$. They are also natural when the target function is well-approximated by a single compositional step with many output channels. While this weight sharing imposes a structural constraint, Sprecher’s theorem guarantees it suffices for universal approximation in single-layer networks, and our experiments suggest it remains effective in deep compositions over the depths explored in Section 8.
Having established the scalability advantages of SNs, we now detail the architecture (Section 2), provide theoretical analysis (Section 6), and present additional empirical demonstrations (Section 8).
@@ 2 Motivation and overview of Sprecher Networks
While MLPs are the workhorse of deep learning, architectures inspired by Kolmogorov–Arnold superposition (KAS) representations offer potential benefits, particularly in interpretability and potentially parameter efficiency for certain function classes. KANs explore one direction by placing learnable functions on edges. Our Sprecher Networks (SNs) explore a different direction, aiming to directly implement Sprecher’s constructive formula within a trainable framework and extend it to deeper architectures.
@@
  • Each functional block (mapping between layers) is organized around a shared monotone spline $\phi(\cdot)$ and a shared general spline $\Phi(\cdot)$, both learnable.
@@
  • Each block includes learnable mixing weights $\lambda_{i}$ (a vector, not a matrix), shared across all output dimensions.
@@
Concretely, each Sprecher block computes a shared-spline shift-and-sum map parameterized by $(\phi^{(\ell)},\Phi^{(\ell)},\lambda^{(\ell)},\eta^{(\ell)})$ and, when enabled, lateral-mixing parameters $(\tau^{(\ell)},\omega^{(\ell)})$ (see Figure 2 and the formal operator definition below). Lateral mixing (if enabled) modifies the pre-activation $s_{q}^{(\ell)}$ (defined below) before $\Phi^{(\ell)}$, whereas residual connections (if enabled) are applied after $\Phi^{(\ell)}$ and do not introduce lateral coupling across $q$. For scalar outputs (summation-head variant), the outputs of the final Sprecher block are aggregated (via summation); for vector outputs, no final summation is applied.
@@
Throughout this paper, we denote Sprecher Network architectures using arrow notation of the form $d_{\mathrm{in}}\to[d_{1},d_{2},\ldots,d_{L}]\to d_{\mathrm{out}}$, where $d_{\mathrm{in}}$ is the input dimension, $d_{1},d_{2},\ldots,d_{L}]$ represents the hidden layer dimensions (widths), and $d_{\mathrm{out}}$ is the final output dimension of the network. For scalar output ($d_{\mathrm{out}}=1$), we typically aggregate the final block’s outputs by summation; alternatively, scalar output can be obtained via the vector-output formulation by including an additional non-summed output block mapping from $d_{L}$ to $1$ (we state explicitly when this variant is used). For vector output ($d_{\mathrm{out}}>1$), an additional non-summed block maps from $d_{L}$ to $d_{\mathrm{out}}$. For example, $2\to[5,3,8]\to 1$ describes a network with 2-dimensional input, three hidden layers of widths 5, 3, and 8 respectively, and a scalar output (implying the final block’s outputs of dimension 8 are summed). $2\to[5,3]\to 4$ describes a network with 2-dimensional input, two hidden layers of widths 5 and 3, and a 4-dimensional vector output (implying an additional output block maps from dimension 3 to 4 without summation). When input or output dimensions are clear from context, we may use the abbreviated notation $d_{1},d_{2},\ldots,d_{L}]$ to focus on the hidden layer structure.
@@ 3.1 Sprecher block structure
  • Monotone inner spline $\phi$: a shared non-decreasing univariate spline (applied coordinate-wise to shifted inputs).
  • Outer spline $\Phi$: a shared univariate spline applied after summation (no monotonicity constraint).
  • Mixing vector $\lambda\in\mathbb{R}^{d_{\mathrm{in}}}$: shared across all $d_{\mathrm{out}}$ outputs (vector weights, not a matrix).
  • Shift $\eta\in\mathbb{R}$: controls the structured input shift $x_{i}\mapsto x_{i}+\eta q$.
  • Optional lateral mixing: $O(d_{\mathrm{out}})$ parameters that couple neighboring channels before applying $\Phi$.
For details on spline parameterizations, enforcing monotonicity of $\phi$, and (optional) adaptive spline domains, see Sections 7 and 7.5.
@@
Concretely, we first define a single Sprecher block as an operator. Let $B^{(\ell)}:\mathbb{R}^{d_{\ell-1}}\to\mathbb{R}^{d_{\ell}}$ denote the $\ell$-th Sprecher block with parameters $\phi^{(\ell)},\Phi^{(\ell)},\eta^{(\ell)},\lambda^{(\ell)}$, a fixed spacing constant $\alpha$, and optionally $\tau^{(\ell)},\omega^{(\ell)}$. Given an input vector $\mathbf{x}=(x_{1},\dots,x_{d_{\mathrm{in}}})\in\mathbb{R}^{d_{\mathrm{in}}}$, the block computes its output vector $\mathbf{h}\in\mathbb{R}^{d_{\mathrm{out}}}$ component-wise as:
$B^{(\ell)}(\mathbf{x})]_{q}=\Phi^{(\ell)}\Biggl(\,s_{q}^{(\ell)}+\tau^{(\ell)}\sum_{j\in\mathcal{N}(q)}\omega_{q,j}^{(\ell)}s_{j}^{(\ell)}\Biggr),$ (2)
where $s_{q}^{(\ell)}=\sum_{i=1}^{d_{\mathrm{in}}}\lambda_{i}^{(\ell)}\phi^{(\ell)}(x_{i}+\eta^{(\ell)}q)+\alpha q$ represents the pre-mixing activation, $\alpha$ is the fixed channel-spacing constant introduced in Section 2, and $\mathcal{N}(q)$ denotes the neighborhood structure:
@@
  • Cyclic: $\mathcal{N}(q)=\{(q+1)\bmod d_{\mathrm{out}}\}$ with a single weight $\omega_{q,j}^{(\ell)}=\omega_{q}^{(\ell)}$
  • Bidirectional: $\mathcal{N}(q)=\{(q-1)\bmod d_{\mathrm{out}},(q+1)\bmod d_{\mathrm{out}}\}$ with separate weights $\omega_{q,(q-1)\bmod d_{\mathrm{out}}}^{(\ell)}$ and $\omega_{q,(q+1)\bmod d_{\mathrm{out}}}^{(\ell)}$
While $\alpha=1$ maintains theoretical fidelity, alternative values may be explored to improve optimization dynamics in deeper networks. Note that $q$ serves dual roles here: it indexes output channels, and it also acts as the fixed one-dimensional coordinate used by the additive shift terms. Although we write $q=0,\ldots,d_{\mathrm{out}}-1$ for notational convenience, in practice one may instead use any fixed grid of $d_{\mathrm{out}}$ coordinates (e.g., a uniform grid on [1,1][-1,1]) to keep shifts well-scaled. (Note: we use 0-based indexing for output channels $q$ to align with the additive shift $\alpha q$; raw inputs use 1-based indices $i=1,\ldots,d_{\mathrm{in}}$ following Sprecher’s original convention, while intermediate layer outputs $\mathbf{h}^{(\ell)}_{r}$ inherit 0-based indexing from their role as output channels of the preceding block.)
@@
Figure 3: Internal structure of a Sprecher block showing spline sharing. Unlike KANs where each edge has a unique learnable spline, a Sprecher block uses only two shared splines: one monotonic $\phi$ applied to all shifted inputs, and one general $\Phi$ applied to all weighted sums. Conceptually the computation is still dense (each output depends on every input through the weighted sum), but no parameters are tied to individual edges: all nonlinearity is shared via the two splines and the shared mixing vector $\lambda$. The mixing weights $\lambda_{i}$ are shared across all output dimensions. Diversity across outputs arises from the index-dependent shifts $\eta q$ (applied before $\phi$) and $\alpha q$ (added to each sum). This extreme parameter sharing yields $O(d_{\mathrm{in}}+G)$ base parameters per block (plus $O(d_{\mathrm{out}})$ when lateral mixing is enabled) versus $O(d_{\mathrm{in}}\cdot d_{\mathrm{out}}\cdot G)$ for KANs.